For vectors U, V, W, show that U.V = V.U and (U + V).W = (U.W) + (V.W)
1
Expert's answer
2020-03-31T10:48:24-0400

Let the vectors U,V and W be

U=Uxi+Uyj+Uzk

V=Vxi+Vyj+Vzk

W=Wxi+Wyj+Wzk

U.V=U*V

=(Uxi+Uyj+Uzk)*(Vxi+Vyj+Vzk)

=UxVx+UyVy+UzVz


V.U=V*U

=(Vxi+Vyj+Vzk)*(Uxi+Uyj+Uzk)

=VxUx+VyUy+VzUz

Thus U.V=V.U


U(V+W)=U.V+U.W

V+W=(Vxi+Vyj+Vzk)+(Wxi+Wyj+Wzk)

=(Vx+Wx)i+(Wy+Wy)j+(Vz+Wz)k


U(V+W)={Uxi+Uyj+Uzk }{(Vx+Wx)i+(Wy+Wy)j+(Vz+Wz)k}


=Ux(Vx+Wx)+Uy(Wy+Wy)+Uz(Vz+Wz)



U.V=VxUx+VyUy+VzUz

U.W=UxWx+UyWy+UzWz

U.V+U.W=Ux(Vx+Wx)+Uy(Wy+Wy)+Uz(Vz+Wz)

Thus U(V+W)=U.V+U.W